OpenStudy (anonymous):

Let f(x) = x2 – 16. Find f–1(x). ± ±

OpenStudy (anonymous):

find the inverse of the function. replace y to x and x to y \[y=x ^{2}-16\] \[x=y ^{2}-16\] Look for new y \[x+16=y ^{2}\] then take square root to look for y alone we have \\[y=\sqrt{x-16}\]

OpenStudy (anonymous):

i mean \[y=\sqrt{x+16}\]

OpenStudy (anonymous):

actually, since \(f(x)=x^2-16\) is not a one to one function it does not have an inverse

OpenStudy (anonymous):

oh yeah my bad =.="

OpenStudy (anonymous):

you could write \(y=\pm\sqrt{x+16}\) but this is not a function precisely because of the \(\pm\)

OpenStudy (anonymous):

if you like you can restrict the domain of \(f\) and then it would have an inverse that is a function. for example you could say "Let \(f(x)=x^2-16\) for \(x\geq 0\) and then the inverse would be \[f^{-1}(x)=\sqrt{x+16}\]